OpenStudy (amenah8):

WILL MEDAL Optimization: If you have 60 meters of fencing and want to enclose a rectangular area up against a long, straight wall, what is the largest area you can enclose?

OpenStudy (ashauntay):

Thus, w=70

OpenStudy (welshfella):

Have you studied calculus? Optimisation problems?

OpenStudy (amenah8):

Yes! We are just learning Optimization now. I just dont know how to start it

OpenStudy (amenah8):

We can say A=lw, but one side already has the wall, so we do not need fencing for one of those 4 sides

OpenStudy (welshfella):

Ok then we need to find the maximum area of rectangle A = lw where l = length and w = width so the 3 sides = 60 so we have the equation l + 2w = 60 ok so far?

OpenStudy (welshfella):



OpenStudy (welshfella):

the area = lw

OpenStudy (amenah8):

oh okay i get it!

OpenStudy (amenah8):

so then how do we estimate the values of the length and width from the 60meters?

OpenStudy (welshfella):

now we need to find an expression for A in terms of one of the variables:- 2w + l = 60 l = 60 - 2w Substituting in the equation A = lw;- A = w(60 - 2w) OK/

OpenStudy (amenah8):

yeah that makes sense.

OpenStudy (amenah8):

should i distribute the w?

OpenStudy (amenah8):

and then get the derivative?

OpenStudy (welshfella):

yes then you find the derivative dA/dw and equate it to zero Then solve for w

OpenStudy (welshfella):

knowing the value of w you use l + 2w = 60 to find the length of the rectangle. Then of course the area required is l * w.

OpenStudy (amenah8):

so: 60w-2w^2 derivative= 60-4x

OpenStudy (amenah8):

i mean 4w

OpenStudy (amenah8):

60-4w=0 -4/60 = w

OpenStudy (welshfella):

w = 60 / 4

OpenStudy (amenah8):

oh right XD

OpenStudy (amenah8):

W= 15 so L= 45

OpenStudy (welshfella):

60 - 4w = 0 4w = 60 remember to put the derivative = zero.

OpenStudy (amenah8):

So Area=675?

OpenStudy (amenah8):

60-4x=0

OpenStudy (welshfella):

L + 2(15) = 60 L = 60 - 2(15) = 60 - 30

OpenStudy (welshfella):

take your time..

OpenStudy (amenah8):

oh right i forgot to plug them back into the equation

OpenStudy (amenah8):

so L=30 and W=15

OpenStudy (amenah8):

And so area=450

OpenStudy (welshfella):

yeah

OpenStudy (amenah8):

This makes much more sense XD

OpenStudy (amenah8):

Thank you so much!

OpenStudy (welshfella):

yw

OpenStudy (amenah8):

and units would be meters?